In a cartesian plane xy.The line which passes through the origin and is perpendicular to the line of equation 5x-3y-1=0 has for equation?


A. 3x-5x=0


B. 3x+5y=0


C. 5x-3y=0


1
Expert's answer
2021-08-13T12:45:52-0400

Slope of the line 5x-3y-1=0 is "m=\frac{5}{3}."

Slope of the perpendicular line must be "-\frac{1}{m}=-\frac{3}{5}."

B. 3x+5y=0.